### v3

# A Generative Approach to Credit Prediction with Learnable Prompts for Multi-scale Temporal Representation Learning

Abstract
Recent industrial credit scoring models remain heavily reliant on manually tuned statistical learning methods. While deep learning offers promising solutions, its effectiveness is often limited by the complexity of financial data, particularly in long-horizon scenarios. In this work, we propose FinLangNet [^1] https://github.com/leiyu0210/FinLangNet, which addresses credit scoring by reframing it as the task of generating multi-scale distributions of a user’s future behavior. Within this framework, tabular data is transformed into sequential representations, enabling the generation of user embeddings across multiple temporal scales. Inspired by the recent success of prompt-based training in Large Language Models (LLMs), FinLangNet also introduces two types of prompts to model and capture user behavior at both the feature-granularity and user-granularity levels. Experimental results demonstrate that FinLangNet outperforms the online XGBoost benchmark, achieving a 7.2% improvement in KS metric performance and a 9.9% reduction in the relative bad debt rate. Furthermore, FinLangNet exhibits superior performance on public UEA archives, underscoring its scalability and adaptability in time series classification tasks.
Machine Learning, Sequential data mining, representation learning
1 Introduction
Credit risk prediction is a cornerstone for financial institutions to devise effective lending policies and informed decisions evaluating the solvency of borrowers (Genovesi et al., 2023). This analytical endeavor is critical in managing and minimizing loan default risks, which is essential for preserving low bad debt levels and mitigating financial losses in the multi-billion dollar credit industry (Cheng et al., 2020). Credit risk models perform binary classifications to discern good from bad customers, improving overdue risk prediction, and effectively managing bad debt while maintaining profitability (Zhao et al., 2023).
In the field of credit risk for industrial scenarios, user data generally comes from multiple sources such as credit reports and product usage behavior data (Lu and Zhang, 2023). For reasons of stability and interpretability, risk control industry basically relies on XGBoost to handle the irregular multisource financial data. The prevalent practice in the industry involves developing derivative features through feature engineering, followed by feature selection and processed with a XGBoost model (Li et al., 2022). Given the inherent noise in multi-source data and the high-dimensional, sparse, and discrete nature of financial behavior data, which often have a relatively high proportion of missing values (Elia et al., 2023), the XGBoost model generally needs to capitalize on its strengths when combined with human prior knowledge (Song et al., 2023). However, this process is time-consuming and costly, and the model struggles to extract meaningful representational information about users. Moreover, XGBoost faces difficulties in tracking user credit profiles over the long term (Niazkar et al., 2024), which poses a substantial challenge for practical downstream tasks.
To overcome these challenges, recent advancements in time-series (Yao et al., 2022), sequential models (Yousefi and Tosarkani, 2022), and graph models (Xue et al., 2024) have fueled exciting endeavors in the field, such as Derisk (Liang et al., 2023). However, these methods typically focus on prediction or classification at a single time point. While the widely adopted XGBoost in the industry demonstrates strong performance and interpretability at individual time points, its effectiveness declines over different time periods. In practical scenarios, it is crucial to model user behavior across various time windows. Therefore, we reformulate this task as a user behavior representation problem across future time periods.
In this work, we proposed a novel FinLangNet framework to handle multi-source and multi-modal data effectively, where non-sequential data representations, which capture the advantages of modeling non-sequential features such as feature interactions and high-dimensional sparse data, are learned through a DeepFM-based (Guo et al., 2017) architecture, and sequential data are processed by the innovative sequential module, which independently captures user representations from each data source and jointly trains them. Moreover, the sequential module incorporates dual prompts at the feature-granularity and user-granularity levels. The outputs are then jointly trained to capture the overall user representation.
The main contributions are summarized as follows:
  1. 1.
    Unlike traditional credit scoring tasks, we reformulate the problem as a generative task. Specifically, we aim to generate future multi-scale distributional representations based on users’ historical behaviors, enabling a more fine-grained and dynamic understanding of credit risks.
  2. 2.
    We propose a novel framework, FinLangNet, to effectively handle multi-source and multi-modal data. In the sequential module, we introduce two classification prompts: A feature-granularity prompt to learn fine-grained feature-level patterns. A user-granularity prompt to capture user-level representations. The head vector of the sequential module serves as the user representation for each data source.
  3. 3.
    It has been deployed in our Behavior Card service, which is used in the operational model during the loan process. It achieves a 9.9% reduction in the relative bad debt rate, proving its practical utility in financial risk assessment. Notably, our sequential module achieves superior performance on the UEA benchmark.
2 Related Work
Credit risk prediction is a crucial task in the financial sector, typically focused on estimating the likelihood of borrower default over a specific time period. Credit scores, such as FICO (Maiden and Maiden, 2024), are widely used evaluation tools (Jensen, 1992; Bücker et al., 2022), generated by algorithms that analyze various user-related data to assess a borrower’s creditworthiness. Extensive research has explored the application of machine learning techniques for credit risk prediction, including decision-tree-based methods like XGBoost (He et al., 2018), graph models such as GraphSAGE (Balmaseda et al., 2023), ChebConv (Liu, 2022), and their combinations (Fein-Ashley et al., 2024). While deep learning methods are often considered to offer enhanced modeling capabilities, existing studies have found that XGBoost generally outperforms deep learning approaches in this domain (Xu et al., 2021).
The processing of irregular, multi-source financial data is a critical challenge in credit risk prediction. Such data are typically presented in tabular form and often involve high-dimensional features, necessitating effective feature selection techniques. Several methods have been proposed to improve performance, including filter methods (Janane et al., 2023), wrapper methods (Ahadzadeh et al., 2023), and embedded methods (Raghu et al., 2023), which enhance both model accuracy (Xu et al., 2024; Ha et al., 2019; Li et al., 2020) and interpretability (Ma et al., 2018; Xu et al., 2021). Deep learning approaches have also been applied to handle these diverse data types (Gorishniy et al., 2021; Borisov et al., 2022), but they do not consistently outperform XGBoost models (Gorishniy et al., 2021) when working with tabular data.
From data perspective, another approach is to treat financial information as sequential data for processing. Structured data based on temporal sequences, such as transaction records or historical behaviors, can be represented in the standard time-series format, which encompass a wide range of methodologies. Models like EDI, DTWI, and DTWD (Bagnall et al., 2018) rely on calculating distances reflective of temporal warping or deviations in time sequences. MLSTM-FCNs (Karim et al., 2019) combine LSTM and CNN layers for feature generation, while WEASEL-MUSE (Schäfer and Leser, 2017) transforms series into symbolic representations. In the domain of credit risk, time-series models include approaches from both statistical and machine learning (El-Qadi et al., 2022), as well as deep learning methods (Forough and Momtazi, 2022; Ala’raj et al., 2021; Wang and Xiao, 2022; Liang et al., 2023) such as LSTM (Yu et al., 2019) and Transformer (Vaswani et al., 2017). However, existing methods are predominantly built on preprocessed, well-structured financial time-series data, which limits their ability to handle irregular, multi-source records. Additionally, most approaches rely on relatively plain model architectures. To overcome these limitations, we propose FinlangNet, a novel approach for credit risk prediction that transforms irregular muti-source financial data into sequential formats. Through a carefully designed prompt mechanism and network modules tailored to different data types, FinlangNet significantly improves credit risk prediction performance in real-world scenarios.
3 Preliminary
Our task is to learn an effective representation of the potential risk in a user’s future credit behavior by leveraging irregular, cross-modal user data. The data we used is derived from real user data of the company’s financial business and is presented in tabular form, comprising three main parts: i) Basic Information, ii) Credit Report and iii) Loan Behavior. We systematically disassembled and reorganized these multi-source data, with each source containing various dimensions, into several distinct parts. For contents with temporal information, we unified them in chronological order, enabling processing in a sequential format.
Figure 1: The dataset consists of two parts: Non-Sequential Data and Sequential Data. Non-Sequential Data includes static user information. Sequential Data represents time-dependent information. Together, they provide a comprehensive view of user characteristics and behaviors over time.
Overall, we categorized the multi-source data into two types: i) non-sequential data, Basic Information $m$, where $m\in\mathbb{R}^{M}$; ii) sequential data, including the Credit Report (which contains Inquiry Record and Account Record) and Loan Behavior information, each comprising various attributes. An overview of the data is presented in Figure 1. More details about datasets are provided in Appendix Table 5. We denote all these sequential data as $z=\{z_{1},z_{2},\cdots,z_{S}\}$, where $z_{s}\in\mathbb{R}^{C\times T}$. The capital $S$ stands for the number of different data sources (In our case, $S$ sources include Inquiry Record and Account Record from Credit Report, and Loan Behavior information), while $s$ represents each source. $C$ is the number of features, reflecting market data, transaction details, financial indicators and other quantitative measurements. $T$ indicates the sequence length of each feature. For data from different sources, $C$ and $T$ take different values.
The input is defined as $X=(m,z)$. The objective is to predict a probability distribution of overdue behavior across different time scales, expressed as: $f_{\theta}:X\rightarrow[0,1]^{L},$ where $L$ represents the number of time scales and each entry in $[0,1]^{L}$ indicates the predicted probability of overdue behavior at a specific time scale. The model is trained under supervision using ground-truth labels, indicating whether overdue behavior occurred at each corresponding time scale. These labels guide the model in learning an accurate probability distribution over future time horizons.
4 Methodology
As Figure 2 illustrates, our framework consists of two parts: a sequential module and a non-sequential module. For the sequential module, we proposed SRG (Sequence Representation Generator) module to process multi-source sequential data. The output from the sequential and non-sequential modules are fused to form a unified shared embedding. Finally, this output is passed through different task to generate representations across different time scales.
Figure 2: FinLangNet Framework Overview. The architecture incorporates two pivotal sub-modules to harness both sequential and non-sequential data effectively. The Sequential Module(SRG) for sequential features. These components are independently trained during the intermediary phase.
4.1 Sequential Module
As shown in Figure 2 within the sequential module, we designed a module called SRG (Sequence Representation Generator) to process sequence data. Each sequential data source $z_{s}\in\mathbb{R}^{C\times T}$ (where $s\in\{1,2,...,S\}$ refers to the index of a data source) undergoes a transformation using a specialized module $f_{\text{SRG}}$, which is responsible for learning temporal dependencies and effectively representing sequential dynamics. The transformation can be expressed as: $O_{\text{ns}_{s}}=f_{\text{SRG}}(z_{s})$, where $O_{\text{ns}_{s}}$ represents the processed output of the $s$-th data source after temporal feature learning.
Considering the issues of data sparsity and model complexity in financial scenarios, we address these problems by applying discretization to sequential data during the data processing stage. Details can be found in Appendix B.1. The tokenizer is a function that converts features in $\mathbb{R}$ into discrete integer tokens in $\mathbb{N}$. For each channel $c$, the tokenizer is defined as: $\mathcal{T}_{c}:\mathbb{R}^{T}\to\mathbb{N}^{T}$
  • Input: A sequential feature vector $w_{c}=(w_{c,1},w_{c,2},\ldots,w_{c,T})\in\mathbb{R}^{T}$ for channel $c$.
  • Output: A sequence of discrete integer tokens $t_{c}=(t_{c,1},t_{c,2},\ldots,t_{c,T})\in\mathbb{N}^{T}$.
To incorporate a global representation inspired by BERT’s [class] token, we prepend a prompt of feature granularity $\widetilde{\phi}_{c}$ to the sequence representation for each channel. For each tokenized sequence $t_{c}=(t_{c,1},t_{c,2},\ldots,t_{c,T})$, the modified sequence is defined as: $t_{c}^{\prime}=(\widetilde{\phi}_{c},t_{c,1},t_{c,2},\ldots,t_{c,T}),$ where $\widetilde{\phi}_{c}\in\mathbb{N}$ is a unique special prompt specifically assigned to represent the entire channel $c$. Thus, the final representation for all channels becomes: $t=\{t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{C}^{\prime}\}$.
After prepending the feature prompt $\widetilde{\phi}_{c}$, we process each modified sequence $t_{c}^{\prime}$ with an embedding layer to transform the discrete tokens into dense representations. Specifically, for each channel $c$, we use an embedding layer $E_{c}:\mathbb{N}\to\mathbb{R}^{d}$, where $d$ is the embedding dimension. This yields the sequence representation: $h_{c}=(h_{c,0},h_{c,1},h_{c,2},\ldots,h_{c,T})\in\mathbb{R}^{(T+1)\times d},$ where: $h_{c,0}=E_{c}(\widetilde{\phi}_{c})$ is the embedding of the prompt $\widetilde{\phi}_{c}$, $h_{c,j}=E_{c}(t_{c,j})$, $\forall j\in\{1,2,\ldots,T\}$, represents the embeddings for the individual tokens of the sequence.
Thus, for all features $c\in\{1,2,\ldots,C\}$, we obtain the representations:$\{h_{1},h_{2},\ldots,h_{C}\},$where each $h_{c}$ is the densely embedded representation of the modified token sequence for feature channel $c$.
Inspired by the recent success of prompt-based training in Large Language Models (LLMs), we aim to use a learnable vector to represent different data sources. Specifically, to construct the final representation $H^{\prime}$, a summary prompt $P_{s}\in\mathbb{R}^{d}$ is concatenated with the aggregated feature $h^{\prime}$.
To compute the global representation $h^{\prime}$ from the features $h_{c}\in\mathbb{R}^{(T+1)\times d}$ along the $C$ components, we perform a weighted sum pooling operation across the $C$ components as follows:
$h^{\prime}=\sum_{c=1}^{C}w_{c}h_{c},\quad h^{\prime}\in\mathbb{R}^{(T+1)\times d},$ (1)
where the weights $w_{c}$ are generated by a weight function $a(h_{c})$, ensuring $\sum_{c=1}^{C}w_{c}=1$.
Next, we concatenate the learnable vector $P_{s}$ (representing the summary prompt) with the global feature $h^{\prime}$ to get the final representation $H^{\prime}$:
$H^{\prime}=[P_{s};h^{\prime}],\quad H^{\prime}\in\mathbb{R}^{(T+2)\times d}.$ (2)
Here, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes the concatenation operation. The resulting representation $H^{\prime}$ combines both the learnable vector $P_{s}$ and the aggregated global features $h^{\prime}$, providing enriched semantic information for downstream tasks.
The process begins with the input representation $H^{\prime}$ and concludes with the final output $O_{\text{ns}_{S}}$. The input $H^{\prime}\in\mathbb{R}^{(T+2)\times d}$ is processed through a Transformer encoder, which consists of multiple layers of self-attention and feed-forward modules. Let $H^{(l)}\in\mathbb{R}^{(T+2)\times D}$ represent the input to the $l$-th layer, where $H^{(0)}=H^{\prime}$, and $D$ denotes the embedding dimension of the Transformer.
Each Transformer layer begins by computing the multi-head attention outputs. The outputs of $k$ attention heads are concatenated and projected back to the dimension $D$ using a learnable projection matrix $W_{O}$:
$H_{\text{attn}}^{(l)}=[\text{head}_{1};\text{head}_{2};\dots;\text{head}_{k}]W_{O}.$ (3)
A residual connection is then combined with the input of the layer, followed by layer normalization to stabilize training:
$H_{\text{out}}^{(l)}=\text{LayerNorm}(H^{(l)}+H_{\text{attn}}^{(l)}).$ (4)
The output is then passed through a two-layer feed-forward network (FFN) for further transformation. The FFN is defined as:
$\text{FFN}(x)=\text{ReLU}(xW_{1}+b_{1})W_{2}+b_{2},$ (5)
where $W_{1},W_{2}\in\mathbb{R}^{D\times D}$ and $b_{1},b_{2}\in\mathbb{R}^{D}$ are learnable parameters of the network.
To retain consistency with previous layers, a second residual connection is applied to the FFN output, followed by layer normalization:
$H^{(l+1)}=\text{LayerNorm}(H_{\text{out}}^{(l)}+\text{FFN}(H_{\text{out}}^{(l)})).$ (6)
This process is repeated for $L$ layers. $H^{(L)}$ represents the final output of the Transformer encoder, where $H^{(L)}\in\mathbb{R}^{(T+2)\times D}$. We extract the first token vector, representing the output of corresponding position $P_{s}$ as the feature expression learned by the current source. The final output $O_{\text{ns}_{S}}$ is computed: $O_{\text{ns}_{S}}=H^{(L)}[0]\in\mathbb{R}^{D}.$
After processing each individual data source through $f_{\text{SRG}}$, the outputs $O_{\text{ns}_{1}},O_{\text{ns}_{2}},\dots,O_{\text{ns}_{S}}$ are finally concatenated along the source dimension to generate a unified representation of the entire multi-source dataset. Formally, the concatenation process can be expressed as:
$O_{\text{ns}}=[O_{\text{ns}_{1}};O_{\text{ns}_{2}};...;O_{\text{ns}_{S}}].$ (7)
The aggregated output $O_{\text{ns}}$ serves as the sequential feature representation of the entire dataset, capturing temporal patterns as well as multi-source dependencies.
Through this approach, the method effectively learns both individual temporal dependencies from each sequential source and cross-source relationships, providing a comprehensive representation for the sequential feature set $z$.
4.2 Non-Sequential Module
As described in Fig 2. within the non-sequential module. Given a set of non-sequential features denoted as $m\in\mathbb{R}^{M}$, we process $m$ using the DeepFM (Guo et al., 2017) model to effectively capture both low-order and high-order feature interactions. We construct a custom dictionary that maps potential feature values or their ranges to unique tokens. Due to the potentially large feature space, we employ a hashing technique to quickly map continuous feature values to integer indices, approximating dense feature representations without storing a full mapping. These integer indices are then passed to the embedding layer of the DeepFM model, where the sparse indices are transformed into dense low-dimensional latent vectors.
The embedding vectors $\{V_{m}\}_{m=1}^{M}$ generated for each feature are passed to the FM component to model pairwise (second-order) interactions between features. This is achieved by computing the inner product $\langle V_{i},V_{j}\rangle$, which measures the interaction strength between feature $i$ and feature $j$. Additionally, the FM component provides a linear modeling of first-order feature importance. The FM output is defined as:
$y_{\text{FM}}=\langle w,m\rangle+\sum_{j_{1}=1}^{M}\sum_{j_{2}=j_{1}+1}^{M}\langle V_{j_{1}},V_{j_{2}}\rangle\cdot m_{j_{1}}\cdot m_{j_{2}},$ (8)
where $w\in\mathbb{R}^{M}$ is the weight for first-order features, and $V_{j}\in\mathbb{R}^{k}$ (where $k$ is user-defined) is the latent vector for feature $j$. Both $w$ and $V_{j}$ are jointly learned during training.
The same embedding vectors are also fed into the deep component, which is a feed-forward neural network designed to learn high-order non-linear feature interactions.
$y_{\text{DNN}}=\text{MLP}(\frac{1}{2}\left[\left(\sum_{i=1}^{M}m_{i_{1}}\mathbf{v}_{i_{1}}\right)^{2}-\sum_{i=1}^{M}\left(m_{i_{1}}^{2}\cdot\mathbf{v}_{i_{1}}^{2}\right)\right]).$ (9)
The final prediction of the DeepFM model combines the outputs of the FM and DNN components:
$O_{m}=\text{MLP}(y_{\text{FM}}+y_{\text{DNN}}).$ (10)
This comprehensive processing effectively leverages both linear and non-linear interactions between non-sequential features, enhancing the model’s ability to capture complex patterns within the static feature set $m$.
Table 1: Performance Metrics for Multiple Models across Multiple Labels. The models in the upper part of the table use tabular data for modeling (internal features are aligned), and the models in the lower part use time series and non-time series features (internal features are also aligned). Other training and evaluation parameters remain consistent.
Label
$y_{1}(\tau=1)$
$y_{2}(\tau=2)$
$y_{3}(\tau=3)$
$y_{4}(\tau=4)$
$y_{5}(\tau=5)$
$y_{6}(\tau=6)$
Avg KS
Metrics
AUC
KS
AUC
KS
AUC
KS
AUC
KS
AUC
KS
AUC
KS
XGBoost(baseline)
72.78
32.85
75.76
37.42
70.89
30.00
73.04
33.18
68.69
26.96
69.70
28.31
31.45
MLP
71.97
31.81
74.76
36.14
69.95
28.70
72.00
31.76
67.59
25.38
68.45
26.55
30.06
GraphSAGE
72.19
32.04
75.09
36.49
70.26
29.09
72.37
32.16
68.03
26.00
68.95
27.25
30.51
ChebConv
72.30
32.11
75.19
36.47
70.39
29.20
72.52
32.22
68.19
26.19
69.16
27.27
30.58
LSTM
72.73
32.86
76.10
38.09
71.01
30.21
73.62
34.33
69.27
27.76
70.98
30.43
32.28
GRU
72.59
32.40
75.68
37.16
70.93
30.05
73.37
33.57
69.06
27.44
70.62
29.75
31.73
Transformer
72.54
32.62
75.95
37.98
70.97
30.12
73.76
34.54
69.30
27.82
71.19
30.67
32.29
Mamba
72.28
32.06
75.66
37.28
70.65
29.67
73.16
33.47
68.79
26.88
70.23
29.02
31.40
TimesNet
72.49
32.54
75.90
37.98
70.83
29.99
73.48
34.15
69.26
27.73
71.05
30.53
32.15
FinLangNet(ours)
73.55
34.08
76.96
39.46
71.92
31.60
74.51
35.67
70.33
29.27
72.12
32.16
33.71
  • $y_{i}$ represents the task label corresponding to the representation of the observation period at different time scales
  • $\tau$ represents different time scales (such as 60 days, 90 days, 180 days, etc.)
4.3 Design of Multi-scale Distribution Representation and Loss Function
4.3.1 Multi-scale Distribution Representation
In our application scenario, we hope to model the representation of users at different time scales in the future so that we can more comprehensively evaluate the credit risk of a user. In our model, the output vector is defined as:
$O=[O_{m};O_{\text{ns}}].$ (11)
Suppose we aim to learn representations at $n$ different scales. The shared representation $O$ is first processed through a multi-layer perceptron and subsequently passed through $n$ separate heads, generating a set of vectors: $A=\{A_{1},A_{2},\dots,A_{n}\}$. Each $A_{i}$ corresponds to the user’s representation at a specific future time scale. To intuitively evaluate the model’s performance, we map each $A_{i}$ to a downstream task $y_{i}$, where $y_{i}$ represents the overdue status of the user at the corresponding time scale, resulting in predictions $\mathbf{y}=\{y_{1},y_{2},\dots,y_{n}\}$.
4.3.2 Loss Function
This design dynamically adjusts the model’s focus on different samples. This loss combines the Mean Squared Error (MSE) and Weighted Logarithmic Loss (WLL), while dynamically adjusting sample-level weights to emphasize harder examples. The total loss function is defined as:
$\mathcal{L}_{\text{total}}=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left[\beta(y^{\prime}_{i}-y_{i})^{2}-(1-\beta)\left(\mathcal{L}_{\text{WLL},i})\right)\right],$ (12)
where the dynamic sample weight $\omega_{i}=\frac{\left(\frac{1}{g_{i}+\epsilon}\right)^{\alpha}}{\sum_{j=1}^{N}\left(\frac{1}{g_{j}+\epsilon}\right)^{\alpha}}$, and the gradient magnitude $g_{i}=\left|\frac{\partial\mathcal{L}_{i}}{\partial y^{\prime}_{i}}\right|.$ Here, $\beta\in[0,1]$ is a hyperparameter controlling the trade-off between MSE and WLL. The Weighted Logarithmic Loss (WLL) is defined as:
$\mathcal{L}_{\text{WLL},i}=$ $-\Big{(}w_{i}^{+}y_{i}\log(y^{\prime}_{i}+\epsilon)$
$+w_{i}^{-}(1-y_{i})\log(1-y^{\prime}_{i}+\epsilon)\Big{)},$ (13)
where $w_{i}^{+}$ and $w_{i}^{-}$ represent the weights for positive and negative samples, and $\epsilon$ is a small constant to prevent numerical instability. Details can be found in Appendix B.2.
5 Experiment
5.1 Experiment Setup
Initially, our experiments are carried out on real business data. We sampled data from over 7 million users, resulting in a dataset of approximately 2 million samples. The dataset contains 6 evaluation tasks, representing overdue status at different time scales. The evaluation indicators of each task are KS (Massey Jr, 1951) and AUC. We compared the online benchmark model and some deep model methods under the same training parameter conditions. Additional details of our dataset are provided in Appendix Table 5. Detailed training parameters can be found in Appendix Table 7.
Furthermore, we have chosen to evaluate the scalability of our framework’s sequential module by analyzing 5 multivariate time series classification (MTSC) tasks sourced from the publicly available UEA archive (Bagnall et al., 2018). These datasets encompass a diverse array of practical applications, including gesture and action recognition, audio analysis, and medical diagnostics through heartbeat monitoring. Our assessment entails a comparison of our FinLangNet with 11 existing MTSC SOTA methods, such as TStamp (Zerveas et al., 2021) and SVP-T (Zuo et al., 2023) . See the Appendix for more details.
5.2 Performance on Credit Prediction
As demonstrated in Table 1, we evaluated the performance of our proposed model across six different time scales, denoted as $\tau$ ranging from 1 to 6. For each corresponding time scale, we labeled the results as $y_{1}$ through $y_{6}$. To benchmark our approach, we compared it against several models, including XGBoost (online baseline models), MLP, and graph-based models such as GraphSAGE and ChebConv. Our model was trained just once and was able to generalize across all six time scales without the need for retraining. Our evaluation was based on key metrics such as AUC and KS. As shown in Table 1, when dealing with tabular data in the risk control domain, XGBoost remains one of the most competitive methods compared to other deep learning approaches. The results clearly demonstrate that our model outperformed all baseline approaches across every individual time scale. On average, compared with online baseline models, our model achieved a 7.2% improvement. This significant improvement highlights the robustness and scalability of our approach.
Moreover, given that our experiment dataset is composed of both sequential and non-sequential data, with the majority being sequential, we conducted additional ablation studies on the sequence modeling component. Specifically, we replaced our sequence module with several other models, including LSTM, GRU, Mamba, Transformer, and TimesNet. The results in Table 1 indicate that our model consistently outperformed all of these alternatives, demonstrating its effectiveness in handling sequential data.
In summary, the experimental results validate the superiority of our model in both average performance and adaptability across multiple time scales. Furthermore, its modularity and efficiency make it highly scalable for various downstream tasks, especially in scenarios with heavy sequential data dependencies.
Table 2: Performance Metrics for Model across Multiple Labels Tested with Various Lengths of Time-Series Data
Label
$y_{1}(\tau=1)$
$y_{2}(\tau=2)$
$y_{3}(\tau=3)$
$y_{4}(\tau=4)$
$y_{5}(\tau=5)$
$y_{6}(\tau=6)$
Avg KS
Metrics
AUC
KS
AUC
KS
AUC
KS
AUC
KS
AUC
KS
AUC
KS
1.00 * Length
73.55
34.08
76.96
39.46
71.92
31.60
74.51
35.67
70.33
29.27
72.12
32.16
33.71
0.50 * Length
73.45
34.00
76.80
39.34
71.80
31.55
74.45
35.61
70.22
29.14
72.07
32.11
33.62
0.25 * Length
73.36
33.82
76.77
39.18
71.72
31.30
74.37
35.54
70.14
29.07
71.99
32.00
33.49
5.3 Ablation Study
5.3.1 Impact of Sequential Data Length
Considering that different users have varying sequence lengths ($T$), it is sometimes necessary to perform sequence truncation or segmentation to handle these variations. To investigate the impact of sequence length on model performance, we conducted ablation experiments where only the length of the sequential data was varied, while keeping non-sequential data constant. As shown in Table 2, the full-length data yielded an average KS value of 33.71%. This decreased marginally to 33.62% at half-length and further to 33.49% at a quarter-length.
Table 3: Influence of Different Modules in Ablation Studies on the $y_{1}(\tau=1)$
Configuration KS AUC
Baseline (No Prompts) 0.3262 0.7254
+ Prompt ($\widetilde{\phi}_{c}$) 0.3279 0.7265
+ Prompt ($P_{s}$) 0.3299 0.7278
+ Both Prompts ($\widetilde{\phi}_{c}$ + $P_{s}$) 0.3408 0.7355
These findings suggest that although there was a reduction in the length of the sequential data, there was only a slight decrease in model performance, which may indicate a certain robustness of the model to changes in sequence length. It also highlights the importance of key information contained within the sequential data for maintaining model performance. Additionally, the results demonstrate the model’s capacity to adapt to variations in data streams that are typical in real-life situations.
5.3.2 Role of Different Module
The second area of our ablation study, as shown in Table 3, aims to investigate the contributions of the two prompts introduced in SRG (our sequence module). Specifically, the first prompt is the feature granularity prompt $\widetilde{\phi}_{c}$, which captures domain-specific feature-level contextual information, while the second prompt is the data-source granularity prompt $P_{s}$, designed to encode interactions and patterns at the data-source level.
Our results show that both prompts contribute positively to model performance, as evidenced by consistent improvements across key evaluation metrics when either prompt is introduced. However, a closer examination reveals that the data-source granularity prompt $P_{s}$ provides a more significant performance gain compared to the feature granularity prompt $\widetilde{\phi}_{c}$. This result suggests that capturing relationships and dependencies at the data-source level is particularly crucial for improving the model’s understanding and utilization of sequential data.
Additionally, we hypothesize that the larger gain from $P_{s}$ is due to its ability to disentangle and organize information from disparate data sources more effectively, improving the model’s ability to generalize and capture global patterns. On the other hand, $\widetilde{\phi}_{c}$ enhances the model’s fine-grained representation at the feature level, complementing $P_{s}$. This synergy between the two prompts explains their combined impact in boosting overall performance (as shown in Table 1), where the integration of both $\widetilde{\phi}_{c}$ and $P_{s}$ results in the best overall performance.
In summary, these findings highlight the importance of incorporating both feature-level and data-source-level contextual information into the model. While each type of prompt contributes independently to the performance enhancement, the greater impact of $P_{s}$ underscores the value of effectively leveraging data-source granularity in sequential modeling. These insights further solidify the advantages of SRG’s design for robust and interpretable sequence prediction tasks.
Figure 3: Model Performance on the UEA archive.
Table 4: Swap Set Analysis on the Performance of Online Baseline Model (XGBoost) and FinLangNet in Actual Business
[…]
(-inf, 478.0] 41.45% 22.36% 14.78% 8.68% 8.00% 35.5% 100.0% 16.7%
(478.0, 560.0] 31.44% 20.42% 14.54% 9.77% 6.40% 20.5% 79.9% 11.9%
(560.0, 634.0] 28.58% 18.89% 13.05% 8.97% 5.70% 14.0% 59.9% 9.1%
(634.0, 730.0] 27.09% 16.60% 11.56% 7.72% 4.69% 8.9% 40.0% 6.6%
(730.0, inf] 27.78% 14.50% 10.35% 6.61% 3.08% 4.2% 20.0% 4.2%
Bad_Rate 38.43% 20.13% 13.04% 7.98% 3.54%
Pass_Rate 100.00% 79.93% 59.81% 39.74% 19.89%
Pass_Bad_Rate 16.67% 11.20% 8.20% 5.76% 3.54%
RBDR 9.90%
  • Yellow_cell - No change (Users who are not replaced under the current decision)
  • Green_cell - Swapping in user (Users whose historical decision is Reject and whose new decision is Approve)
  • Orange_cell - Swapping out user (Users whose historical decision is Approve and whose new decision is Reject)
  • Bad_Rate - The overdue rate of users in the corresponding binning
  • Pass_Rate - The proportion of users who pass the corresponding binning decision
  • Pass_Bad_Rate - The overdue rate of users under the corresponding binning decision
  • RBDR - Relatively bad decline rate. RBDR = (Pass_Bad_Rate(Baseline) - Pass_Bad_Rate(Ours)) / (Pass_Bad_Rate(Baseline))
5.4 Scalability
We also evaluate the scalability of our method to demonstrate its benefits in time series classification tasks. The accuracy comparison results, illustrated in Figure 3, showcase the superior performance of our methods across the majority of public datasets. These results underscore the versatility of our model, highlighting its effectiveness across various real-world domains. Please refer to the detailed results presented in Appendix Table 8.
6 Online A/B Test
In practical business applications, the predicted probability output by the model is transformed into a credit score using a predefined fixed-parameter function. A higher predicted probability of delinquency corresponds to a lower credit score, and vice versa. During the decision-making phase, user scores are partitioned into distinct binning intervals through equal-frequency binning. This process facilitates the determination of a cutoff threshold, where applications with scores above the cutoff are approved, while those below it are rejected. Typically, there exists a positive correlation between the approval rate and the bad debt rate, meaning that an increase in the approval rate often results in a corresponding rise in the bad debt rate.
Although technical metrics such as AUC and KS provide valuable insights into model performance, the ultimate goal of the model is its integration into downstream risk control strategies. In the deployment of credit risk models, Swap Set Analysis (Kritzinger and van Vuuren, 2019) is commonly employed to assess performance. This analysis involves “swapping out” a subset of customer accounts—typically those classified as high-risk (bad accounts)—and “swapping in” a new set of accounts, generally consisting of lower-risk (good) customers. This approach provides a more intuitive understanding of the model’s impact on credit decision-making and overall portfolio risk.
Table 4 presents the Swap Set Analysis comparing our model with the baseline in a real business scenario. The table’s rows and columns represent the distribution of credit scores assigned by the two models to the same batch of customers, with bin boundaries determined using equal-frequency binning into five groups. By mapping the scores of both models into a 5×5=25 grid, we obtain a comparative view of their performance. The values in the table indicate the actual overdue rate (bad rate) within each bin.
Beyond the table, we calculate the bad rate for each bin, the cumulative pass rate, and the bad rate of passed samples (i.e., the actual overdue rate of customers approved under the given cutoff). In practical business applications, the bad rate of passed samples is the key performance metric. Swap Set Analysis allows us to compare the new and baseline models under the same approval rate to observe differences in the bad rate of passed samples.
For illustration, we analyze a cutoff decision boundary at a 60% pass rate, meaning that the top 60% of customers are approved, while the bottom 40% are rejected. Under this setting, the baseline model exhibits a bad rate of 9.1% for the approved population, whereas our new model achieves a lower bad rate of 8.2% under the same pass rate. This indicates an absolute risk reduction of 0.9%, corresponding to a relative improvement of 9.9% over the baseline. As a result, our model effectively mitigates nearly 10% of potential asset losses, delivering substantial economic benefits to the credit business.
7 Conclusion
In this work, we propose an innovative framework FinLangNet to solve practical credit risk business problems. We treat credit scoring as a generative task, predicting multi-scale distributions of users’ future behavior. The design of FinLangNet incorporates a DeepFM-based architecture for non-sequential data and the innovative SRG module for sequential data. In SRG module, we introduce dual prompts at the feature and user-granularity levels, allowing for fine-grained feature learning and holistic user representation. Experimental results demonstrate FinLangNet’s effectiveness, achieving a 7.2% improvement in KS metric and a 9.9% reduction in the relative bad debt rate compared with online baseline. Furthermore, FinLangNet achieves superior performance on public UEA archives, showcasing its scalability and adaptability in time-series classification tasks.
Appendix A Data
A.1 The Credit Dataset Details
We sampled real data from over 700,000 active users to participate in the modeling. The sample size level is about 2 million. The assembled datasets consist of three primary sources illustrating different dimensions of consumer financial behavior. The visualization presented in Figure 4, demonstrates a data visualization approach for a subset of our dataset. We randomly selected 256 individuals and focused on four distinct features per individual. From these features, only 10% of their lengths were extracted to perform a focused analysis. The more details is provided in Appendix Table 5.
Figure 4: Feature Sub-Section Heatmap of Selected users.
Delving into the datasets themselves, the Basic Information dataset aggregates a compilation of over 700,000 entries, with each entry endowed with 11 attributes. These attributes relay a combination of demographic insights and rudimentary financial information on the subjects. Concurrently, the Credit Report dataset is divided into two intricate sub-datasets: Firstly, the Inquiry Feature sub-dataset encapsulates the trajectory of credit inquiries with more than 700,000 entries, each chronicled through a sequence of 120 timestamped records, and each record is intricately detailed with 5 distinctive features. Secondly, the Account Feature sub-dataset expands this vista by supplying over 700,000 entries each backed by 200 time-based observations, and an elaborate array of 29 different features for every observation, presenting a fulsome view of individual account behaviors over time. Complementing these, the Loan Behaviour dataset meticulously records transactional data concerning loans for the same sprawling number of subjects, parsed into 26 discrete time points, each defined by 241 unique attributes.
Cumulatively, these datasets form an extensive resource for analyzing consumer credit patterns, with a breadth of information ranging from personal identifiable information to intricate details of credit inquiries, account management, and loan repayment behaviors.In the intricate realm of industrial scenarios, the employment of meticulously processed datasets is quintessential. As portrayed in the table 5, the data employed in this study is the fruition of a complex transformation process, hailing from disparate data sources. Each segment has undergone a tailored procedure to distill the information into a usable format for analysis and model training. The foundation of our dataset is the Basic Information segment, where we extract key personal and financial details from relevant text sources. This extraction focuses on gathering the most pertinent attributes, ensuring a solid base for our dataset.
Within the Credit Report database, we divide the data into Inquiry and Account segments, each offering unique insights into the credit reporting process. Through sequence alignment, we achieve a unified and chronological view of the data, which facilitates a more nuanced understanding of credit behavior over time. The Loan Behaviour data is managed through time series analysis, which allows us to navigate its temporal complexities. This approach precisely tracks the evolution of loan transactions, capturing the authentic rhythm of financial movements. Together, these datasets present a comprehensive and nuanced view of the industrial dataset landscape, primed for detailed exploration and advanced modeling. Each segment’s specialized handling enriches the overall data pool, providing a robust foundation for rigorous analysis.
Table 5: Dataset descriptions.
Data Source Channel Sequential Length Dataset Size
Description
Basic Information
11 None ($\sim$3000000, 0, $\sim$600000)
Personal Information
Credit Report: Inquiry
5 120 ($\sim$3000000, 0, $\sim$600000)
Credit Records
Credit Report: Account
29 200 ($\sim$3000000, 0, $\sim$600000)
Credit Records
Loan Behaviour
241 26 ($\sim$3000000, 0, $\sim$600000)
Borrowing or Repayment Behavior
A.2 The UEA Dataset Details
For the purpose of testing and validation, we utilized the core module on the UEA archive. Nonetheless, we undertook a random selection encompassing all 5 types of multivariate time series datasets available within the UEA archive. Details of these datasets are shown in Table 6.
Table 6: UEA Classification Dataset Details
Dataset Train Cases Test Cases Dimensions Length Classes AtrialFibrillation(AF) 15 15 2 640 3 BasicMotions(BM) 40 40 4 100 4 EthanolConcentration(EC) 261 263 3 1751 4 Heartbeat(HB) 204 105 61 495 2 SelfRegulationSCP2(SCP2) 200 180 7 1152 2
Appendix B Method
B.1 Data Processing
For the input data, the features of each channel $c\in\{1,2,\ldots,C\}$ are classified into two types: continuous feature channel set $\mathcal{C}_{\text{cont}}$ and categorical feature channel set $\mathcal{C}_{\text{cat}}$, where
$|\mathcal{C}_{\text{cont}}|=C_{\text{cont}},\quad|\mathcal{C}_{\text{cat}}|=C_{\text{cat}},$
and the total number satisfies $C_{\text{cont}}+C_{\text{cat}}=C$.
For continuous feature channels $c\in\mathcal{C}_{\text{cont}}$, the input is a temporal feature vector:
$W_{c}=(w_{c,1},w_{c,2},\ldots,w_{c,T})\in\mathbb{R}^{T}.$
These continuous features are discretized through a quantization function $\mathcal{Q}_{c}:\mathbb{R}\to\mathbb{N}$, resulting in a discrete integer token sequence:
$t_{c}=(t_{c,1},t_{c,2},\ldots,t_{c,T})\in\mathbb{N}^{T},\quad t_{cj}=Q_{c}(w_{cj}),\;\forall j\in\{1,\ldots,T\}.$
For categorical feature channels $c\in\mathcal{C}_{\text{cat}}$, the input is a temporal categorical feature vector:
$W_{c}=(w_{c,1},w_{c,2},\ldots,w_{c,T}),\quad w_{cj}\in\mathcal{X}_{c},$
where $\mathcal{X}_{c}$ represents the set of categorical features for channel $c$. These categorical features are discretized via a hash encoding function $\mathcal{H}_{c}:\mathcal{X}_{c}\to\{1,2,\ldots,H_{c}\}$, producing the token sequence:
$t_{c}=(t_{c,1},t_{c,2},\ldots,t_{,cT})\in\mathbb{N}^{T},\quad t_{cj}=\mathcal{H}_{c}(w_{cj}),\;\forall j\in\{1,\ldots,T\}.$
Finally, the discretized outputs for all channels are unified into a single representation as discrete integer token sequences:
$t_{c}=(t_{c1},t_{c2},\ldots,t_{cT})\in\mathbb{N}^{T},\quad\forall c\in\{1,2,\ldots,C\}.$
For $c\in\mathcal{C}_{\text{cont}}$:
$t_{cj}=Q_{c}(w_{cj});$
For $c\in\mathcal{C}_{\text{cat}}$:
$t_{cj}=\mathcal{H}_{c}(w_{cj}).$
This unified discretization approach ensures that all features, regardless of type, are represented in a consistent format, enabling further processing and modeling.
B.2 Loss Function
To address the issue of class imbalance in our binary classification tasks, we propose a modified loss function that incorporates both a reweighting scheme and dynamic adjustment mechanisms. Specifically, at each iteration, the loss weights are recalculated based on the gradient norms, allowing the model to focus more on harder examples. We denote by $\omega_{r}$ the weight for the $r$-th loss component, adjusted using a tuning parameter $\alpha$.
Our total loss function $\mathcal{L}_{\text{total}}$ combines the weighted Dice-BCE loss and the Focal Tversky loss as follows:
$\mathcal{L}_{\text{total}}=\omega_{1}\mathcal{L}_{\text{Dice-BCE}}+\omega_{2}\mathcal{L}_{\text{Focal-Tversky}},$ (14)
where $\omega_{1}$ and $\omega_{2}$ are the dynamic weights for each loss component.
The weighted Dice-BCE loss $\mathcal{L}_{\text{Dice-BCE}}$ is defined as:
$\mathcal{L}_{\text{Dice-BCE}}=\lambda\mathcal{L}_{\text{Dice}}+(1-\lambda)\mathcal{L}_{\text{BCE}},$ (15)
where $\lambda\in[0,1]$ balances the contribution between the Dice loss and the Binary Cross-Entropy (BCE) loss.
The Dice loss $\mathcal{L}_{\text{Dice}}$ is calculated by:
$\mathcal{L}_{\text{Dice}}=1-\frac{2\sum_{i=1}^{N}p_{i}y_{i}+\epsilon}{\sum_{i=1}^{N}p_{i}+\sum_{i=1}^{N}y_{i}+\epsilon},$ (16)
and the BCE loss $\mathcal{L}_{\text{BCE}}$ is given by:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right],$ (17)
where $p_{i}\in[0,1]$ is the predicted probability for the $i$-th sample, $y_{i}\in\{0,1\}$ is the corresponding binary label, $N$ is the total number of samples, and $\epsilon$ is a small constant to prevent division by zero.
The Focal Tversky loss $\mathcal{L}_{\text{Focal-Tversky}}$ is defined as:
$\mathcal{L}_{\text{Focal-Tversky}}=\left(1-\text{Tversky Index}\right)^{\gamma},$ (18)
where $\gamma\geq 1$ is the focal parameter that emphasizes harder-to-classify examples.
The Tversky Index is calculated by:
Tversky $\text{Tversky Index}=\frac{\sum_{i=1}^{N}p_{i}y_{i}+\epsilon}{\sum_{i=1}^{N}p_{i}y_{i}+\alpha\sum_{i=1}^{N}p_{i}(1-y_{i})+\beta\sum_{i=1}^{N}(1-p_{i})y_{i}+\epsilon},$ (19)
where $\alpha$ and $\beta$ are constants that control the penalties for false positives and false negatives, respectively.
The dynamic weights $\omega_{r}$ are recalculated at each iteration based on the gradient norms $G_{r}$ of each loss component:
$\omega_{r}=\frac{\left(\frac{1}{G_{r}+\epsilon}\right)^{\alpha}}{\sum_{k}\left(\frac{1}{G_{k}+\epsilon}\right)^{\alpha}},$ (20)
where
$G_{r}=\left\|\nabla_{\theta}\mathcal{L}_{r}\right\|.$ (21)
$\nabla_{\theta}\mathcal{L}_{r}$ is the gradient of the $r$-th loss component with respect to the model parameters $\theta$, and $\alpha$ is a tuning parameter that adjusts the sensitivity of the weight updates.
By incorporating this dynamic reweighting scheme, the model adaptively emphasizes loss components that contribute more to the overall gradient, effectively handling class imbalance. The combined loss function $\mathcal{L}_{\text{total}}$ facilitates improved convergence and performance by balancing the influence of each loss component according to the current training dynamics.
This design allows the predicted probabilities $p_{i}$ (ranging from 0 to 1) to be effectively compared with the binary labels $y_{i}$ in the loss computations, optimizing the model’s performance on the binary classification tasks across different downstream applications.
The proposed loss function builds on the principles of Dice-BCE loss and Focal Tversky loss, combining their emphasis on focusing on harder examples with dynamic adjustment of sample weights. Below, we detail the formulation of the loss function step by step.
The Weighted Logarithmic Loss is defined as:
$\mathcal{L}_{\text{WLL},i}=$ $-\Big{(}w_{i}^{+}y_{i}\log(y^{\prime}_{i}+\epsilon)+w_{i}^{-}(1-y_{i})\log(1-y^{\prime}_{i}+\epsilon)\Big{)},$ (22)
where: - $y_{i}$ is the ground truth label for sample $i$, and $y^{\prime}_{i}$ is the predicted value. - $w_{i}^{+}$ and $w_{i}^{-}$ are weights for positive and negative examples, respectively, used to handle class imbalance. - $\epsilon>0$ is a small constant for numerical stability.
This logarithmic loss penalizes predictions more harshly when they deviate significantly from the true labels.
To prioritize harder examples, dynamic weights are calculated for each sample based on the gradient magnitude of the loss with respect to the prediction. The sample weight $\omega_{i}$ is defined as:
$\omega_{i}=\frac{\left(\frac{1}{g_{i}+\epsilon}\right)^{\alpha}}{\sum_{j=1}^{n}\left(\frac{1}{g_{j}+\epsilon}\right)^{\alpha}}.$ (23)
$g_{i}=\left|\frac{\partial\mathcal{L}_{i}}{\partial y^{\prime}_{i}}\right|$ is the gradient magnitude for the $i$-th sample. $\alpha$ is a tunable exponent that controls the emphasis on harder examples (higher $\alpha$ increases their importance). $\epsilon>0$ ensures numerical stability.
This weight assigns higher importance to samples with larger gradient magnitudes, focusing the training process on harder examples.
The final total loss function combines the Mean Squared Error (MSE) term and the Weighted Logarithmic Loss (WLL) term, controlled by a trade-off parameter $\beta\in[0,1]$. For each sample $i$, the loss is expressed as:
$\mathcal{L}_{i}=\omega_{i}\left[\beta(y^{\prime}_{i}-y_{i})^{2}-(1-\beta)\mathcal{L}_{\text{WLL},i}\right].$ (24)
The MSE term, $(y^{\prime}_{i}-y_{i})^{2}$, penalizes the squared error between the prediction and ground truth. The WLL term, $\mathcal{L}_{\text{WLL},i}$, introduces a logarithmic penalty for classification errors. - $\beta$ balances the trade-off between MSE and WLL.
The total loss over all $n$ samples is:
$\mathcal{L}_{\text{total}}=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left[\beta(y^{\prime}_{i}-y_{i})^{2}-(1-\beta)\mathcal{L}_{\text{WLL},i}\right].$ (25)
Substituting the definition of WLL into the equation, the total loss becomes:
$\mathcal{L}_{\text{total}}=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\Big{[}$ $\beta(y^{\prime}_{i}-y_{i})^{2}-(1-\beta)\Big{(}w_{i}^{+}y_{i}\log(y^{\prime}_{i}+\epsilon)+w_{i}^{-}(1-y_{i})\log(1-y^{\prime}_{i}+\epsilon)\Big{)}\Big{]}.$ (26)
Dynamic Weighting ($\omega_{i}$): Focuses on harder examples by assigning larger weights to samples with higher gradient magnitudes. Trade-off Hyperparameter ($\beta$): Controls the balance between the MSE term (for regression-like behavior) and the WLL term (for robust classification). By integrating MSE and WLL dynamically, the approach benefits from both regression accuracy and log-scale penalties, while adapting to the varying difficulty of examples. This loss formulation retains the focus on harder examples, similar to Focal Tversky and Dice-BCE, while unifying classification and regression objectives in a single framework.
Appendix C Experiments
C.1 Training Details and Evaluation Metric
The integration of these metrics, along with the Kolmogorov-Smirnov statistic, furnishes a multi-faceted evaluation framework for the risk assessment model. The Kolmogorov-Smirnov statistic, by measuring the maximum divergence in cumulative distributions between positive and negative classes, directly addresses the model’s discriminative power. Similarly, the AUC provides a measure of the model’s ability to differentiate between classes over the entire range of classification thresholds. Table 7 shows the key hyperparameters used in the training of the risk assessment model and their set values.
Parameter Value Description
Optimizer AdamW Optimizer with weight decay for better generalization.
Learning Rate 0.0005 Step size for parameter updates.
Betas (0.9, 0.999) Decay rates for first and second moment estimates in Adam.
Epsilon 1e-08 Smoothing term to avoid division by zero.
Weight Decay 0.01 Regularization to prevent overfitting.
Balancing Coefficients $\alpha$ 0.5 Balances different loss components.
Balancing Coefficients $\beta$ 0.5 Balances regression (MSE) and classification (WLL).
Balancing Coefficients $\gamma$ 1.0 Focuses on hard-to-learn samples.
Epochs 12 Number of complete passes through the training dataset.
Table 7: Training Parameters and Descriptions.
C.2 Case Study
In a case study, we analyzed two models: XGBoost (XGB) and FinLangNet. The XGB utilized manually engineered features informed by domain expertise and relied on a single feature for making predictions. FinLangNet, however, constructed sequential features from a subset of the original features using advanced algorithms to extract insights without preconceived notions, and it broadened its scope by applying six different labels to enrich the feature set.
A fair comparison required focusing on the common label $y_{1}(\tau=1)$. The evaluation of model performance was conducted over a one-month future testing window. In the context of risk control, the threshold marking a ”high risk” case is essential. As seen in Figure 5, the XGB achieves high Recall at low thresholds, capturing most positive cases but with reduced Precision, leading to a high false positive rate. Conversely, FinLangNet displays more balanced performance, offering high Precision and slightly lower Recall at low thresholds, resulting in fewer misclassifications.
Figure 5: Model Performance Comparison at Various Thresholds on the $y_{1}(\tau=1)$.
Figure 6: Comparison of Predictive and the $y_{1}(\tau=1)$ Distributions for XGBoost and FinLangNet.
Our examination of the $y_{1}(\tau=1)$ label as illustrated in Figure 6 revealed that FinLangNet outperforms the XGBoost model in identifying users at high risk of default. Even though it sometimes incorrectly classifies low-risk users, FinLangNet was found to be more adaptable and robust against imbalanced labels, ensuring it retains its superior performance in skewed datasets.
C.3 Multivariate Time Series Classification Task
C.3.1 Baselines Details
For classification, we refer to the SOTA methods: EDI, DTWI, and DTWD [Bagnall et al., 2018] rely on calculating distances reflective of temporal warping or deviations in time sequences. MLSTM-FCNs [Karim et al., 2019] combine LSTM and CNN layers for feature generation, while WEASEL-MUSE [Schäfer and Leser, 2017] transforms series into symbolic representations. Scalable Representation Learning (SRL) [Franceschi et al., 2019] builds representations via negative sampling and encoders, TapNet [Zhang et al., 2020] is a recent model with an attentional prototype learning in its deep learning-based network, ShapeNet [Li et al., 2021] discovers shapelets through clustering in a unified space, and Rocket [Dempster et al., 2020] employ random convolutional kernels for rapid feature extraction from univariate series. These methods like TStamp [Zerveas et al., 2021] and SVP-T [Zuo et al., 2023] ollectively elevate the precision and understanding of temporal data analysis.
C.3.2 Results
Impressively, even bereft of the full suite of FinLangNet capabilities, the core module’s performance was observed to either mirror or exceed the benchmarks established by the currently trending models in the domain. The specifics of our performance metrics and the accuracy scores for these datasets are meticulously detailed in Table 8.
It is important to note that two most recently SOTA models in Multivariate Time Series Classification, such as TimesNet [Wu et al., 2022] and ShapeFormer [Le et al., 2024]. We tested the performance of TimesNet on our dataset, as shown in Table 1 . Regarding the shapeFormer, although it performs exceptionally well on the UEA archive, it could not be validated in our setting. This limitation is due to the model’s dependency on the Shapelet module, which requires the extraction of discriminative sub-sequences from the entire dataset in advance. Our scenario, characterized by an ultra-large-scale data stream, makes the implementation of this approach impractical. Therefore, we were unable to verify the method. Nonetheless, we have compared the performances of TimesNet, ShapeFormer, and FinLangNet across different datasets in Table 8 and 9.
Table 8: Accuracies on Five Datasets of the UEA Archive
EDI DTWI DTWD MLSTM-FCNs WEASEL+MUSE SRL TapNet ShapeNet Rocket TStamp SVP-T FinLangNet AF 0.267 0.267 0.220 0.267 0.333 0.133 0.333 0.400 0.249 0.200 0.400 0.400 BM 0.676 1.000 0.975 0.950 1.000 1.000 1.000 1.000 0.990 0.975 1.000 1.000 EC 0.293 0.304 0.323 0.373 0.430 0.236 0.323 0.312 0.447 0.337 0.331 0.414 HB 0.619 0.658 0.717 0.663 0.727 0.737 0.751 0.756 0.718 0.712 0.790 0.795 SCP2 0.483 0.533 0.539 0.472 0.460 0.556 0.550 0.578 0.514 0.589 0.600 0.600
Table 9: Comparison with Two Most Recently Advanced TSC Methods on Public UEA and Our Dataset.
TimesNet ShapeFormer FinLangNet AF $\diamond$ N/A 0.600 0.400 BM $\diamond$ N/A 1.000 1.000 EC $\diamond$ 0.357 0.378 0.414 HB $\diamond$ 0.780 0.800 0.795 SCP2 $\diamond$ 0.572 0.633 0.600 Our Dataset $\ast$ 0.322 N/A 0.337
  • $\ast$ Using KS as metric; $\diamond$ Using Acc as metric.

### v4

@@
 While deep learning offers promising solutions, its effectiveness is often limited by the complexity of financial data, particularly in long-horizon scenarios.
-In this work, we propose FinLangNet [^1] https://github.com/leiyu0210/FinLangNet, which addresses credit scoring by reframing it as the task of generating multi-scale distributions of a user’s future behavior.
+In this work, we propose FinLangNet [^1] https://github.com/didiglobal-fintech-credit-risk/FinLangNet, which addresses credit scoring by reframing it as the task of generating multi-scale distributions of a user’s future behavior.
 Within this framework, tabular data is transformed into sequential representations, enabling the generation of user embeddings across multiple temporal scales.
@@
 Through a carefully designed prompt mechanism and network modules tailored to different data types, FinlangNet significantly improves credit risk prediction performance in real-world scenarios.
+Meanwhile, we also investigated several LLM-based approaches.
+In real-world applications, general-purpose large language models often fall short in meeting specific business needs.
+For example, although CALM(Feng et al., 2023) is a financial-domain LLM, it is not tailored for credit scoring tasks.
+CALM outputs binary classifications (e.g., Good/Bad), while our model, FinLangNet, produces continuous delinquency probabilities, which can be directly mapped to credit scores.
+These scores are crucial for making fine-grained decisions such as customer acceptance or rejection based on cutoff thresholds.
+Moreover, model size and output stability are also critical considerations in production environments, where large models like CALM may not perform reliably or efficiently.
 3 Preliminary
@@
 This comprehensive processing effectively leverages both linear and non-linear interactions between non-sequential features, enhancing the model’s ability to capture complex patterns within the static feature set $m$.
-Table 1: Performance Metrics for Multiple Models across Multiple Labels.
-The models in the upper part of the table use tabular data for modeling (internal features are aligned), and the models in the lower part use time series and non-time series features (internal features are also aligned).
-Other training and evaluation parameters remain consistent.
+Table 1: Performance comparison across multiple models and labels.
+The upper part lists models using tabular features, while the lower part includes models combining sequential and non-sequential features.
+All models share the same training and evaluation settings.
 Label
@@
 31.45
-MLP
+MLP (Goodfellow et al., 2016)
 71.97
@@
 30.06
-GraphSAGE
+GraphSAGE (Hamilton et al., 2017)
 72.19
@@
 30.51
-ChebConv
-72.30
-32.11
-75.19
-36.47
-70.39
-29.20
-72.52
-32.22
-68.19
-26.19
-69.16
-27.27
-30.58
-LSTM
-72.73
-32.86
-76.10
-38.09
-71.01
-30.21
-73.62
-34.33
-69.27
-27.76
-70.98
-30.43
-32.28
-GRU
+GRU (Cho et al., 2014)
 72.59
@@
 31.73
-Transformer
+Transformer (Vaswani et al., 2017)
 72.54
@@
 32.29
-Mamba
+Mamba (Gu and Dao, 2023)
 72.28
@@
 31.40
-TimesNet
+TimesNet (Wu et al., 2023)
 72.49
@@
 32.15
+DIN (Zhou et al., 2018)
+73.16
+33.63
+76.51
+38.98
+71.56
+31.16
+74.13
+35.19
+70.03
+28.89
+71.84
+31.67
+33.25
+DIEN (Zhou et al., 2019)
+73.45
+33.83
+76.91
+39.33
+71.72
+31.30
+74.42
+35.33
+70.03
+28.79
+71.86
+31.58
+33.36
+BST (Chen et al., 2019)
+73.28
+33.76
+76.80
+39.35
+71.62
+31.23
+74.37
+35.52
+69.95
+28.78
+71.89
+31.75
+33.40
 FinLangNet(ours)
@@
 where $w_{i}^{+}$ and $w_{i}^{-}$ represent the weights for positive and negative samples, and $\epsilon$ is a small constant to prevent numerical instability. Details can be found in Appendix B.2.
-5 Experiment
-5.1 Experiment Setup
-Initially, our experiments are carried out on real business data.
-We sampled data from over 7 million users, resulting in a dataset of approximately 2 million samples.
-The dataset contains 6 evaluation tasks, representing overdue status at different time scales.
-The evaluation indicators of each task are KS (Massey Jr, 1951) and AUC.
-We compared the online benchmark model and some deep model methods under the same training parameter conditions.
-Additional details of our dataset are provided in Appendix Table 5.
-Detailed training parameters can be found in Appendix Table 7.
-Furthermore, we have chosen to evaluate the scalability of our framework’s sequential module by analyzing 5 multivariate time series classification (MTSC) tasks sourced from the publicly available UEA archive (Bagnall et al., 2018).
-These datasets encompass a diverse array of practical applications, including gesture and action recognition, audio analysis, and medical diagnostics through heartbeat monitoring.
-Our assessment entails a comparison of our FinLangNet with 11 existing MTSC SOTA methods, such as TStamp (Zerveas et al., 2021) and SVP-T (Zuo et al., 2023) .
-See the Appendix for more details.
-5.2 Performance on Credit Prediction
-As demonstrated in Table 1, we evaluated the performance of our proposed model across six different time scales, denoted as $\tau$ ranging from 1 to 6.
-For each corresponding time scale, we labeled the results as $y_{1}$ through $y_{6}$.
-To benchmark our approach, we compared it against several models, including XGBoost (online baseline models), MLP, and graph-based models such as GraphSAGE and ChebConv.
-Our model was trained just once and was able to generalize across all six time scales without the need for retraining.
-Our evaluation was based on key metrics such as AUC and KS.
-As shown in Table 1, when dealing with tabular data in the risk control domain, XGBoost remains one of the most competitive methods compared to other deep learning approaches.
-The results clearly demonstrate that our model outperformed all baseline approaches across every individual time scale.
-On average, compared with online baseline models, our model achieved a 7.2% improvement.
-This significant improvement highlights the robustness and scalability of our approach.
-Moreover, given that our experiment dataset is composed of both sequential and non-sequential data, with the majority being sequential, we conducted additional ablation studies on the sequence modeling component.
-Specifically, we replaced our sequence module with several other models, including LSTM, GRU, Mamba, Transformer, and TimesNet.
-The results in Table 1 indicate that our model consistently outperformed all of these alternatives, demonstrating its effectiveness in handling sequential data.
-In summary, the experimental results validate the superiority of our model in both average performance and adaptability across multiple time scales.
-Furthermore, its modularity and efficiency make it highly scalable for various downstream tasks, especially in scenarios with heavy sequential data dependencies.
 Table 2: Performance Metrics for Model across Multiple Labels Tested with Various Lengths of Time-Series Data
@@
 33.49
+5 Experiment
+5.1 Experiment Setup
+Initially, our experiments are carried out on real business data.
+We sampled data from over 7 million users.
+The dataset contains 6 evaluation tasks, representing overdue status at different time scales.
+The evaluation indicators of each task are KS (Massey Jr, 1951) and AUC.
+We compared the online benchmark model and some deep model methods under the same training parameter conditions.
+Additional details of our dataset are provided in Appendix Table 5.
+Detailed training parameters can be found in Appendix Table 7.
+Furthermore, we have chosen to evaluate the scalability of our framework’s sequential module by analyzing 5 multivariate time series classification (MTSC) tasks sourced from the publicly available UEA archive (Bagnall et al., 2018).
+These datasets encompass a diverse array of practical applications, including gesture and action recognition, audio analysis, and medical diagnostics through heartbeat monitoring.
+Our assessment entails a comparison of our FinLangNet with 11 existing MTSC SOTA methods, such as TStamp (Zerveas et al., 2021) and SVP-T (Zuo et al., 2023) .
+See the Appendix for more details.
+5.2 Performance on Credit Prediction
+As illustrated in Table 1, we evaluate model performance across six temporal scales, denoted as $\tau$ ranging from 1 to 6, with prediction results labeled as $y_{1}$ through $y_{6}$.
+Each task is assessed using AUC and KS as evaluation metrics.
+In the upper part of the table, we compare traditional tabular-based models.
+XGBoost, the current online baseline, exhibits strong performance across most tasks and remains highly competitive in risk control scenarios.
+However, our proposed FinLangNet consistently outperforms XGBoost and other deep learning baselines on all six tasks, achieving an average improvement of 7.2% in key metrics.
+This demonstrates both the robustness and scalability of our framework in handling complex business data.
+The lower section of Table 1 focuses on models tailored for sequential inputs, including GRU (Cho et al., 2014), Transformer (Vaswani et al., 2017), Mamba (Gu and Dao, 2023), TimesNet (Wu et al., 2023), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019) and BST (Chen et al., 2019).
+All models are trained under identical hyperparameter settings.
+Even in this setting, FinLangNet achieves superior performance with negligible additional computational cost.
+We attribute the superior performance of our model to its enhanced ability to generalize across heterogeneous channels and data sources, a limitation commonly observed in standard sequential models that often struggle in real-world scenarios despite performing well on academic datasets.
+To address this, we introduce two tailored prompts that facilitate cross-source information integration.
+Furthermore, we reformulate credit risk modeling as a multi-scale representation learning task, enabling comprehensive user risk profiling across the full credit lifecycle.
+This multi-scale design promotes mutual learning across time horizons, contributing to consistently optimal performance across all tasks.
 5.3 Ablation Study
 5.3.1 Impact of Sequential Data Length
-Considering that different users have varying sequence lengths ($T$), it is sometimes necessary to perform sequence truncation or segmentation to handle these variations.
-To investigate the impact of sequence length on model performance, we conducted ablation experiments where only the length of the sequential data was varied, while keeping non-sequential data constant.
-As shown in Table 2, the full-length data yielded an average KS value of 33.71%.
-This decreased marginally to 33.62% at half-length and further to 33.49% at a quarter-length.
-Table 3: Influence of Different Modules in Ablation Studies on the $y_{1}(\tau=1)$
-Configuration KS AUC
-Baseline (No Prompts) 0.3262 0.7254
-+ Prompt ($\widetilde{\phi}_{c}$) 0.3279 0.7265
-+ Prompt ($P_{s}$) 0.3299 0.7278
-+ Both Prompts ($\widetilde{\phi}_{c}$ + $P_{s}$) 0.3408 0.7355
+To assess the effect of sequence length on model performance, we conducted a set of ablation experiments in which only the length of the sequential inputs was varied, while non-sequential features were held fixed.
+As reported in Table 2, shortening the sequence led to only marginal performance degradation, suggesting that the model exhibits robustness to sequence length variation.
+This also underscores the presence of critical information in early parts of the sequence and highlights the model’s ability to generalize under conditions of incomplete or truncated data.
+Table 3: Influence of Different Modules on task $y_{1}(\tau=1)$
+Configuration KS(%) AUC(%)
+Baseline (No Prompts) 32.62 72.54
++ Prompt ($\widetilde{\phi}_{c}$) 32.79 72.65
++ Prompt ($P_{s}$) 32.99 72.78
++ Both Prompts ($\widetilde{\phi}_{c}$ + $P_{s}$) 34.08 73.55
+w/o Sequential model ($O_{\text{ns}}$) 12.53 58.62
+w/o Non-sequential model ($O_{m}$) 32.75 72.66
 These findings suggest that although there was a reduction in the length of the sequential data, there was only a slight decrease in model performance, which may indicate a certain robustness of the model to changes in sequence length.
@@
 5.3.2 Role of Different Module
-The second area of our ablation study, as shown in Table 3, aims to investigate the contributions of the two prompts introduced in SRG (our sequence module).
-Specifically, the first prompt is the feature granularity prompt $\widetilde{\phi}_{c}$, which captures domain-specific feature-level contextual information, while the second prompt is the data-source granularity prompt $P_{s}$, designed to encode interactions and patterns at the data-source level.
-Our results show that both prompts contribute positively to model performance, as evidenced by consistent improvements across key evaluation metrics when either prompt is introduced.
-However, a closer examination reveals that the data-source granularity prompt $P_{s}$ provides a more significant performance gain compared to the feature granularity prompt $\widetilde{\phi}_{c}$.
-This result suggests that capturing relationships and dependencies at the data-source level is particularly crucial for improving the model’s understanding and utilization of sequential data.
+We conducted ablation studies to evaluate the contribution of different components in our model, as summarized in Table 3.
+First, we assess the impact of the two prompts in the SRG module: the feature-granularity prompt $\widetilde{\phi}_{c}$ and the data-source prompt $P_{s}$.
+Both prompts individually improve performance over the no-prompt baseline, with $P_{s}$ yielding a larger gain, indicating that modeling cross-source interactions is more critical than capturing fine-grained feature context.
 Additionally, we hypothesize that the larger gain from $P_{s}$ is due to its ability to disentangle and organize information from disparate data sources more effectively, improving the model’s ability to generalize and capture global patterns.
@@
 This synergy between the two prompts explains their combined impact in boosting overall performance (as shown in Table 1), where the integration of both $\widetilde{\phi}_{c}$ and $P_{s}$ results in the best overall performance.
-In summary, these findings highlight the importance of incorporating both feature-level and data-source-level contextual information into the model.
-While each type of prompt contributes independently to the performance enhancement, the greater impact of $P_{s}$ underscores the value of effectively leveraging data-source granularity in sequential modeling.
-These insights further solidify the advantages of SRG’s design for robust and interpretable sequence prediction tasks.
+Furthermore, we examine the role of the sequential and non-sequential modules.
+Removing the sequential module results in a substantial performance drop, confirming its critical role in capturing temporal dynamics.
+In contrast, removing the non-sequential module leads to only a slight decline, implying that sequential information carries the dominant signal for our task.
+These results validate the effectiveness of our dual-prompt strategy and modular design, highlighting the importance of incorporating both hierarchical prompts and multi-modal input streams for robust credit risk modeling.
 Figure 3: Model Performance on the UEA archive.
@@
 As a result, our model effectively mitigates nearly 10% of potential asset losses, delivering substantial economic benefits to the credit business.
+For deployment, the model is converted to ONNX format and served on an L20 GPU cluster.
+Given the low-frequency nature of credit applications, the system operates under moderate QPS ($\sim 100$) and achieves inference latency within 100 ms, meeting real-time business requirements.
+To ensure transparency and trustworthiness, we incorporate a lightweight interpretability framework based on the LIME (Local Interpretable Model-agnostic Explanations) (Ribeiro et al., 2016) method.
+LIME perturbs the input—by masking or shuffling elements in the sequence—and learns a local surrogate model to identify the most influential features driving predictions.
+This allows us to interpret the behavior of our sequential model in a locally faithful manner.
+We apply this mechanism across different data sources to rank feature importance, which serves multiple practical purposes: (i) guiding feature selection during model development, (ii) supporting issue diagnosis after deployment, and (iii) monitoring feature drift in production.
 7 Conclusion